OpenStudy (anonymous):

Which of the following forms a proportion with 8/20?.

OpenStudy (anonymous):

\[\frac{ 10 }{ 4 } \] It Has Too Be One Of These

OpenStudy (anonymous):

\[\frac{ 6 }{ 15 }\]

OpenStudy (anonymous):

\[\frac{ 10 }{ 16 } \frac{ 4 }{ 16 }\]